OpenStudy (anonymous):

Emma purchased a refrigerator for $1,450. It depreciates about 0.8% each year. What is the value of the refrigerator after five years?

OpenStudy (anonymous):

8% or \(.8\%\)?

OpenStudy (anonymous):

if it is \(.8\%\) then after 5 years it would be worth \[1450\times (.992)^5\]

OpenStudy (anonymous):

.8%^ your so funny you have answer like all my questions alredy

OpenStudy (anonymous):

1392.92

OpenStudy (anonymous):

it is \(.992\) since \(.8\%=0.008\) and \(1-0.008=.992\)

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

Do you want to help me with my last question of the day?lol Elliot has deposited $723 in a savings account that earns interest at a rate of 2.9% compounded twice a year. What will the account balance be in 23 years?

OpenStudy (anonymous):

\[723\times (1+\frac{.029}{2})^{46}\]

OpenStudy (anonymous):

Your a gift from the gods thank you=3

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=723%281%2B.029%2F2%29^46 about $1401.95 they are easy once you know the formula do you have the formulas for these?

OpenStudy (anonymous):

Yeah the words are just what trows me off=D

OpenStudy (anonymous):

ok well it sounds like you are done in any case

OpenStudy (anonymous):

bye thansk again

OpenStudy (anonymous):

yw again